OpenStudy (anonymous):

I need help solving Quadratic Equations. Comment if you know how to solve them so I can give you the problems.

OpenStudy (precal):

many ways to solve a quadratic: Quadratic formula (always works) Factoring (does not always work especially if the quadratic is prime) complete the square (too tedious) square root property but it has to be set up a certain way graphing method, look at the graph to see the number of solutions

OpenStudy (anonymous):

Solve: 3y2 - 48 = 0

OpenStudy (anonymous):

using the quadratic formula attached , we have the form ax^2 + bx + c = 0 .. here we have a = 3 , b = 0 , c = -48 , substitute into the formula and you will get the result y = 4 or y = -4

OpenStudy (anonymous):

What about this one: 5x2 - 45 = 0

OpenStudy (anonymous):

you do the same steps using the quadratic formula , here we have a = 5 , b = 0 , c = -45 so by substituting in the quadratic formula you get x = 3 or x = -3

OpenStudy (anonymous):

 What about this one?